Finch Training Services (FTS) provides instruction on the use of computer software for the employees of its corporate clients. It offers courses in the clients’ offices on the clients’ equipment. The only major expense FTS incurs is instructor salaries; it pays instructors $5,200 per course taught. FTS recently agreed to offer a course of instruction to the employees of Novak Incorporated at a price of $470 per student. Novak estimated that 20 students would attend the course.

Base your answers on the preceding information.

The instructor has offered to teach the course for a percentage of tuition fees. Specifically, she wants $250 per person attending the class. Assume that the tuition fee remains at $470 per student.

  1. Is the cost of instruction a fixed or a variable cost?

  2. Determine the profit, assuming that 20 students take the course.

  3. Determine the profit, assuming a 10 percent increase in enrollment (i.e., enrollment increases to 22 students). What is the percentage change in profitability?

  4. Determine the profit, assuming a 10 percent decrease in enrollment (i.e., enrollment decreases to 18 students). What is the percentage change in profitability?
Answer #1

F. Cost of instruction is variable cost as it is related to per person cost available in the class.

G.Determination of Profit under g,h,i conditions

g. Base h 10%increase i 10%decrease
Particulars 20 Students 22 Students 18 Students
tuition fees @ $470 per student 9400 10340 8460
Less;-Instructors salaries 5200 5200 5200
Profit 4200 5140 3260

g. Percentage change in Profitability by assuming 10% increase in enrollment = (5140 - 4200) x 100/4200 = 22.38%

h. Percentage change in Profitability by assuming 10% decrease in enrollment = (4200-3260) x 100/4200 = 22.38%

Note;- As per the instructions given in the question " Base your answers on the preceding information". That's  why only instructor's salaries has been taken as an Expense for Finch Training Services While calculating Profit.